OpenStudy (anonymous):

If x = y-3 and y=z^2, what is x in terms of z? Answer: a. z^-6 b. z^6 c. z^-2/3 d. z^5 e. z^3/2 keep getting z^2-3, which isn't one of the answers. I have searched the web for an answer, but haven't found one yet.

OpenStudy (anonymous):

Ok, so you want an equation for x in terms of z. That means that you want the only variables in your function to be x and z. Do you have any idea of how you could do this? (Hint: If you have an equation for a variable, you could plug in that equation in for any instance of the variable in another equation.)

OpenStudy (loser66):

do you have the original problem, if you have a scanner, please, scan and post it. I don't see anything wrong with Twis7ed 's stuff

OpenStudy (anonymous):

Here, you go.

OpenStudy (loser66):

ok, I guess there is some mistypo there. yours might be x = y^3 not y -3

OpenStudy (loser66):

@Twis7ed what do you think? are you with me?

OpenStudy (anonymous):

Yeah, might be y^3 or y^-3 or something like that.

OpenStudy (loser66):

thanks for being with me. @dondean don't worry, I am pretty sure about the mistake from your question.

OpenStudy (anonymous):

OK I will it report to the website, because I couldn't make heads or tails from it.

OpenStudy (anonymous):

I think you are correct @Twis7ed and @Loser66. The answer would be 1/z^6, ie z^(-6). If x=y^(-3)

OpenStudy (anonymous):

Yep :)